OpenStudy (anonymous):

A baseball player hits a ball. The ball leaves the bat with an initial upward velocity of 64 feet per second. How many seconds will it take for the ball to hit the ground? Use the formula in which the height above ground after t seconds is given by the quadratic function h = –16t2 + vt, where h represents the height of the object in feet, and v represents the object's initial velocity in feet per second. 4 seconds 8 seconds 0 seconds 1 of a second 16

OpenStudy (amistre64):

how high off the ground is the ball when it hits the ground?

OpenStudy (amistre64):

if you know how to find the "axis of symmetry" of the quadratic equation; then the time it takes to go up is the same amount of time it takes to come down .... so double the x value

OpenStudy (anonymous):

64 feet high

OpenStudy (amistre64):

its 64 feet high when it hits the ground?

OpenStudy (amistre64):

lets assume for the sake of sanity, that its 0 feet high when it hits the ground \[ h = –16t^2 + vt\]\[ 0 = –16t^2 + vt\] and solve for t \[ 0 = t(–16t+v)\] so t = 0 or t = v/16